OpenStudy (anonymous):

What multiplies to get -6 and adds up to 3

Nnesha (nnesha):

factor -6 and 3

Nnesha (nnesha):

what two number you should multiply to get -6 ? and what two number you should multiply to get 3 ?? what number are divisible by 3?

OpenStudy (fibonaccichick666):

so list the factors of 6. Is the first thing Nnesha wants you to do.

OpenStudy (godlovesme):

@snwjwk WELCOME TO OS :D

OpenStudy (fibonaccichick666):

can you do that @snwjwk ?

OpenStudy (anonymous):

-6*1=-6 2*-3=-6 Adds up to 3 2+1=3

OpenStudy (fibonaccichick666):

ah, but you have 2+-3 that =-1

OpenStudy (fibonaccichick666):

keep going with the factoring

Nnesha (nnesha):

okay nice -6 times 1 = -6 2 times -3 = -6 which pair of two number give u the positive 3 ??

OpenStudy (fibonaccichick666):

maybe mess with which values are negative and which are positive

Nnesha (nnesha):

^yeah play with signs

OpenStudy (anonymous):

I'm stuck and clue less

OpenStudy (fibonaccichick666):

There isn't actually an answer

Nnesha (nnesha):

do you have an equation ?

OpenStudy (fibonaccichick666):

it's a trick question :/ but in general this is your algorithm: list the factors of |-6| until you would repeat a number.\[1*6\\2*3\] Now, if your other number is positive, the smaller numbers must be negative. If the number is negative, the big numbers must win and be negative. So here we have negative little numbers. \[-1*6\\-2*3\] So you add those and are like, none of those work! so we cannot get a solution to this question

OpenStudy (anonymous):

Yeah identify any cvs lies for which expression is undefined X^2-x-2 ---------- 3x-6

OpenStudy (anonymous):

Values

OpenStudy (fibonaccichick666):

oh, use the quadratic formula for that

OpenStudy (fibonaccichick666):

don't try and factor it normally, that only works if it factors into integers

OpenStudy (anonymous):

So it would be no solution

OpenStudy (fibonaccichick666):

no, you just use the quadratic formula to factor it

OpenStudy (fibonaccichick666):

do you know it?

OpenStudy (anonymous):

No

OpenStudy (fibonaccichick666):

ohohoohohoohohohohohoh

OpenStudy (fibonaccichick666):

wait a second.

OpenStudy (anonymous):

K

OpenStudy (fibonaccichick666):

do you have this? \[\frac{x^2-x-2}{3x-6}\]

OpenStudy (anonymous):

Yes

OpenStudy (fibonaccichick666):

ohhkkk, so that just means the denominator =0

OpenStudy (fibonaccichick666):

we cannot divide by zero right?

OpenStudy (fibonaccichick666):

so if it does, we go crazy and can't have that

OpenStudy (anonymous):

It doesn't say

OpenStudy (fibonaccichick666):

^that is fact

OpenStudy (fibonaccichick666):

so set the denominator =0 and solve for x, can you give that a try?

OpenStudy (anonymous):

I don't have to factor the top -2 *-1=-2 -2+-1=-1

OpenStudy (fibonaccichick666):

nope

OpenStudy (anonymous):

Okay

OpenStudy (fibonaccichick666):

it's only when the denominator equals zero that you have a problem

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

What times -25 and adds up 1 Wouldn't that just factor 5?

OpenStudy (fibonaccichick666):

nah just solve 3x-6=0

OpenStudy (fibonaccichick666):

what's x =? That's your answer